2.Consider a point-point link 50 km in length. At what bandwidth would propagation delay (at a speed of 2 x 108 m/s) equal transmit delay for 100-byte packets? What about 512-byte packets?
Answer #1

So, the propagation delay = d / s

where d = distance between the two points

s = speed of propagation

And, the transmit delay = (L / B)

where, L = size of data packet

B = bandwidth

(a) In this problem, we have d = 50 km = 5 * 104 m

s = 2 * 108 m/s

L = 100 bytes = 100 * 8 bits = 800 bits

We have to find B for which transmit delay is equal to propagation delay.

So, here is the solved equation.

So, the required bandwidth is 3.2 Mbps.

(b) Here, we have

d = 50 km = 5 * 104 m

s = 2 * 108 m/s

L = 512 bytes = 512 * 8 bits = 4096 bits.

Here is the solved equation.

The reuired bandwidth is 16.384 Mbps.

Let me know if you have any doubts in the comments. Please upvote if the answer helped you.